Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name jroche
Log entry time 08:02:57 on October 13, 2010
Entry number 334599
keyword=owl shift summary
October 12, 2010, Owl shift summary
Crew:
J. Roche
Y. Zhihong
F. Itard
Timeline:
-
00:00: start with epics gateway crash, called MCC, epics was restored
around 12:15 . We have the sieve slit and carbons optic foils in
position. The replay doesn't show the foils...
1:00 Because we have some doubts about the sieve pisition and we need
statistic on with sieve out, we decide to remove the sieve.
1:20: carbons foils, no sieve : now we see the foils !
1:30: Eric asks us to take Compton data.
2:00: taking some runs with S0 in the trigger for scintilator timing
checkout.
2:36: Putting the sieve back in position with carbon foils. This is just
to check that we still cannot make up the foils. We still cannot of
course. I had had some doubts about the sieve slit position.
3:41: MCC wants to do a calibration of their MPS bcm. I did cross check
with htem that this is not the full BCM calibration that we are waiting
for since a while.
5:15: changing target to LH2 15cm, turning off the raster, sive is in
position. DAQ rate=22 Hz.
6:56: MCC takes the beam away for a QE measurement
6:58: beam is back
Run list:
- 7230: sieve slit on, multifoil carbon, 50 uA, pnom=3.327,
angle=27.494, 135k evt
- 7231: sieve slit on, multifoil carbon, 50 uA, pnom=3.327,
angle=27.494, 96k evt
Sieve out
- 7232: sieve slit OUT, multifoil carbon, 50 uA, pnom=3.327,
angle=27.494, 200k evt
- 7233, 7234: junk, playing out with prescale.
- 7235: sieve slit OUT, multifoil carbon, 50 uA, pnom=3.327,
angle=27.494, 200k evt, s0 in the trigger ps2=25
- 7236: sieve slit OUT, multifoil carbon, 50 uA, pnom=3.327,
angle=27.494, 200k evt, s0 in the trigger ps2=50
Sieve in
- 7237: sieve slit in, multifoil carbon, 50 uA, pnom=3.327,
angle=27.494, 200k evt,
- 7238: sieve slit in, LH2, 4.5 uA, pnom=3.327, angle=27.494
- 7239: sieve slit in, LH2, 4.5 uA, pnom=3.327, angle=27.494